Assume the​following:

i. The public holds no currency.

ii. The ratio of reserves to deposits

​(theta

​)

is

0.08

.

iii. The demand for money is given by

Upper M Superscript d Baseline equals $Y left parenthesis 0.87 minus 2.1 i right parenthesis

.

If the monetary base is

​$83

billion and nominal income is

​$5.7

​trillion,the equilibrium interest rate will be

nothing

​%.

​(Round

your response to two decimal

places​.)
Answer #1

Assume the​following:

i. The public holds no currency so C-D ratio is 0.

ii. The ratio of reserves to deposits ​(theta​) is 0.08 so R-D ratio is 0.08

iii. The demand for money is given by Md = Y(0.87 - 2.1i)

Monetary base is ​$83 billion which implies that money supply is monetary base x money multiplier or 83 x 1/0.08. This gives money supply = $1037.50 billion

Now nominal income is ​$5.7 ​trillion or $5700 billion. Use money demand = money supply relation

Y(0.87 - 2.1i) = 1037.50

5700(0.87 - 2.1i) = 1037.50

2.1i = 0.687982

i = 0.3276 or 32.76%.

This gives nominal rate of interest i = 32.76%